1.1) You lend a friend RM20,000 for a year. At the end of the year your friend agrees to pay you RM21,100. The interest rate on this loan is _______

Working Calculation:


1.2) You lend a friend RM20,000 for a year at an annual interest rate of 5%. At the end of the year your friend must pay you ________ in interest.

Working Calculation:


1.3) You agree to lend ________ to a friend for a year at an annual interest rate of 10%. At the end of the year your friend pays you RM600 in interest.

Working Calculation:








1
Expert's answer
2020-07-29T10:58:23-0400

1.1) The interest rate on this loan is 5.5%

Working calculation:

"A=P(1+r)^{t}"

"RM21,100=RM20,000(1+r)^{1}"


"\frac{RM21,100}{RM20,000}=\frac{RM20,000}{RM20,000}(1+r)"


"1.055=1+r"

"r=1.055-1"

"r=0.055\times100"

"r=5.5\%"


1.2) The interest amount is RM 1,000

Working calculation:

"A=P(1+r)^{1}-P"

"RM20,000(1+0.05)-RM20,000"

"(RM20,000\times1.05)-RM20,000"

"RM21,000-RM20,000"


"Interest=RM1,000"


1.3) The amount lend will be RM 6,000

Working calculation:

"Interest=A-P"

"If 10\%=RM600=Interest"

"P=100\%=?"

"Therefore P=\frac{100\%}{10\%}\times RM600"


"P=RM6,000"


Test to confirm the answer:

"A=P(1+r)^{1}-P"

"RM6,000(1+0.1)^{1}-RM6,000"

"(RM6,000\times1.1)-RM6,000"

"RM6,600-RM6,000"


"Interest\; amount=RM600"